# 4K4DGen: Panoramic 4D Generation at 4K Resolution

### v2

@@ 3.2 Inconsistent Perspective Animation
Large-scale pre-trained 2D models have demonstrated impressive generative capabilities in producing images and videos, leveraging massive multi-modal training data sourced from the Internet. However, collecting high-quality 4D training data is non-trivial, and no existing 4D dataset matches the scale of those available for images and videos. Therefore, our approach aims to utilize the capabilities of video generative models to produce consistent panoramic 360-degree videos, which are then elevated to 4D. Nonetheless, the availability of panoramic videos is significantly more limited compared to planar perspective videos. Consequently, mainstream image-to-video (I2V) animation techniques may not perform optimally on panoramic formats, and the resolution of the videos remains constrained, as illustrated in Fig. 6 (b) and Tab. 2. Alternatively, the animator can be applied to perspective images. However, this method results in inconsistencies among different projected views, as illustrated in Fig. 6 (c).
@@ 3.3 Consistent Panoramic Animation
Limited by the scarcity of 4D training data in panoramic format, and given that large diffusion models are primarily trained on planar perspective videos, directly applying 2D perspective denoisers poses challenges due to inconsistent motion between different perspective views. This limitation has motivated us to develop a panoramic video generator that utilizes priors from image-to-video (I2V) animators, which are predominantly trained on planar perspective image-video pairs. Consequently, given a static input panorama, we animate it into a panoramic video, as demonstrated in the "Animating Phase" section of Fig. 3.
@@
To generate panoramic video from a static panorama, we adopt the denoise-in-latent-space schema utilized in latent diffusion models an2023latent ; blattmann2023stable ; dai2023animateanything . Initially, a noisy latent code is sampled, which is then progressively denoised using the DDPM procedure ho2020denoising . Throughout each step of this denoising process, the static input image serves as a conditioning term. Finally, the clean latent code is decoded by a pre-trained VAE decoder to render the video in pixel space. In 4K4DGen, unlike the method for generating perspective planar videos, both the latent code and the input static panorama are represented on spheres. We start with the initial panoramic latent code $S^{T}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$, where $L$ denotes the number of video frames and $c$ the channels per frame. A novel Panoramic Denoiser then generates the clean panoramic latent code $S^{0}$, conditioned on the static input panorama $I\in{}^{H\times W}$. Subsequently, the equirectangular projection, as introduced in Sec. 3.1, projects the clean panoramic latent code into the matrix-like latent code $Z^{0}\in{}^{h\times w\times L\times c}$, with $h$ and $w$ representing the resolution of the latent code. Each $k^{\rm{th}}$ video frame $I^{k}$ in pixel space is decoded by the pre-trained VAE decoder as $I^{k}=\mathcal{D}(Z^{0}[:,:,k,:])$.
@@
We employ a pre-trained perspective video generative model dai2023animateanything to develop our Panoramic Denoiser. This video generator takes a perspective image $i\in{}^{p_{H}\times p_{W}\times c}$ and an initial latent code $z^{T}\in{}^{p_{h}\times p_{w}\times(L\times c)}$ as inputs. It then progressively denoises the latent code $z^{T}$ to a noise-free state $z^{0}$ using a denoising function $z^{t-1}=\Phi(z^{t},i)$, where $p_{h},p_{w}$ are the resolutions of the latent code, $p_{H},p_{W}$ are the dimensions of the conditioning image, $c$ is the number of channels, and $L$ represents the video length. Our goal is to transform the initial noisy latent code $S^{T}$ into the clean panoramic latent code $S^{0}$, ensuring that each perspective view is appropriately animated while maintaining global consistency. The underlying intuition is that if each perspective view undergoes its respective denoising process, the perspective video will feature meaningful animation. Moreover, if two perspective views overlap, they will align with each other bar2023multidiffusion ; lugmayr2022repaint to produce a seamless global animation. Given a static input panorama $I$ and an initial spherical latent code $S^{0}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$, we progressively remove noise employing a project-and-fuse procedure at each denoising step. Specifically, the spherical latent code at the $t^{\rm{th}}$ denoising step, $S^{t}:\mathbb{S}^{2}\rightarrow{}^{L\times C}$, is projected into multiple perspective latent codes $\mathcal{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots,z^{t}_{n}\}$, where each $z^{t}_{k}=\gamma(S^{t},\boldsymbol{d}_{k})\in{}^{p_{h}\times p_{w}\times(L\times c)}$ represents the $k^{\rm{th}}$ perspective latent code projected in the equirectangular format detailed in Sec. 3.1. Each perspective latent code is then denoised by one step using a pre-trained perspective denoiser, denoted as $z_{k}^{t-1}=\Phi(z^{t}_{k},i_{k})$, where $i_{k}=\gamma(I,\boldsymbol{d}_{k})\in{}^{p_{H}\times p_{W}\times c}$ is the perspective conditioning image projected from the panorama $I$. Subsequently, we optimize the spherical latent code $S^{t-1}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$ at step $t-1$ by fusing all the denoised perspective latent codes ${z_{k}^{t-1}}$. Formally, the denoising procedure at step $t$, denoted as $S^{t-1}=\Psi(S^{t},I)$, encompasses the following operations:
$\Psi\left(\mathcal{S}^{t},I\right)=\operatornamewithlimits{argmin}_{\mathcal{S}}\mathbb{E}_{\boldsymbol{d}\in\mathbb{S}^{2}}\|\gamma(\mathcal{S},\boldsymbol{d})-\Phi\left(\gamma(\mathcal{S}^{t},\boldsymbol{d}),\gamma(I,\boldsymbol{d})\right)\|.$ (3)
3.4 Panoramic 4D Lifting
@@
We define the panoramic video as $V=\{I^{1},I^{2},\ldots,I^{L}\}$, consisting of $L$ frames. The video is divided into overlapping perspective videos $\{v_{0},v_{1},\ldots,v_{n}\}$, each captured from specific camera directions $\{\boldsymbol{d}_{1},\ldots,\boldsymbol{d}_{n}\}$, collectively encompassing the entire span of the panoramic video $V$. Subsequently, we estimate the geometry of the 4D scene by fusing the depth maps through spatial-temporal geometry alignment. Following this, we describe our methodology for 4D representation and the subsequent rendering procedure.
@@
To transition from 2D video to 3D space, we utilize a monocular depth estimator ranftl2021vision , inspired by advancements in zhou2024dreamscene360 , to estimate the scene’s geometric structure. Nonetheless, depth maps generated for each frame and perspective might lack spatial and temporal consistency. To address this, we implement Spatial-Temporal Geometry Alignment using a pre-trained depth estimator $\Theta:{}^{h\times w\times 3}\rightarrow{}^{h\times w}$, applied to perspective images. Our objective is to amalgamate $n$ perspective depth maps $D^{K}_{i}=\Theta(\gamma(I^{k},\boldsymbol{d}_{i}))$ into a cohesive panoramic depth map $D^{k}$ for each frame $I^{k}$, ensuring spatial and temporal continuity. We express these depth maps as a spherical representation ${\mathcal{S}_{D}^{1},\ldots,\mathcal{S}_{D}^{L}}$. For enhanced optimization, we assign $n$ scale factors $\alpha_{i}^{k}\in\real$ and shifting parameters $\beta_{i}^{k}\in{}^{h\times w}$ to each perspective depth map. The comprehensive depth map $D^{k}$ is then optimized jointly with these parameters $\alpha$ and $\beta$. The formal objective is structured as follows:
@@
Figure 4: Visualization Results for 4K4DGen. Displayed are the input panorama (Pano. RGB), the optimized global geometry (Pano. Depth), and the final rendered results. For each panorama, we illustrate the rendered video and depth frames from three timestamps ($t=0$, $t=0.5$, $t=1$), captured from two overlapping views (View1, View2). 4K4DGen effectively generates 4D scenes that are both spatially and temporally consistent.
@@
We represent and render the dynamic scene using $T$ sets of 3D Gaussians. Each set, corresponding to a specific timestamp $t$, is denoted as $G_{t}=\{\left(\boldsymbol{p}_{t}^{i},\boldsymbol{q}_{t}^{i},\boldsymbol{s}_{t}^{i},\boldsymbol{c}_{t}^{i},o_{t}^{i}\right)|i=1,\ldots,n\}$. This definition aligns with the methods described in bahmani20234d , which also provides a fast rasterizer for rendering images based on these Gaussian sets and given camera parameters. Consistent with Sec. 3.1, while the camera intrinsics remain fixed, we parameterize the camera extrinsics through a position $\boldsymbol{p}\in{}^{3}$ and an orientation $\boldsymbol{d}\in\mathbb{S}^{2}$. The training process is structured in two stages: initially, we directly supervise the 3D Gaussians using the panoramic videos. Let $\mathcal{R}(G,\boldsymbol{p},\boldsymbol{d})$ represent the rasterized image from Gaussian set $G$, utilizing camera extrinsics $\boldsymbol{p}=0$ and camera direction $\boldsymbol{d}$. Let $I_{t}$ denote the $t^{\rm{th}}$ frame of the panoramic video. We optimize the $t^{\rm{th}}$ Gaussian set $G_{t}$ using the following objective:
@@
where the RGB supervision term $\mathcal{L}_{\rm{rgb}}=\lambda\mathcal{L}_{1}+(1-\lambda)\mathcal{L}_{\rm{SSIM}}$ is the same as 3D-GS kerbl20233d , and the temporal regularize term $\mathcal{L}_{\rm{temporal}}$ written as:
$\mathcal{L}_{\rm{temporal}}=\sum\limits_{i=1}^{n}\|\mathcal{R}(G_{t},\boldsymbol{0},\boldsymbol{d}_{i})-\mathcal{R}(G_{t-1},\boldsymbol{0},\boldsymbol{d}_{i}))\|$ (6)
In the second stage, we adopt the distillation loss and geometric regularization used in zhou2024dreamscene360 , the distillation loss is defined as follows: $\mathcal{L}_{sem}=1-\operatorname{cos}\left\langle\operatorname{CLS}(\mathcal{R}(G_{t},\boldsymbol{0},\boldsymbol{d}_{i})),\operatorname{CLS}(\mathcal{R}(G_{t},\boldsymbol{\delta}_{p},\boldsymbol{d}_{i}))\right\rangle$, where $\boldsymbol{\delta}_{p}\in[-\alpha,\alpha]^{3}$ is the disturbing vector, $\operatorname{CLS}(\cdot)$ the feature extractor such as DINO oquab2023dinov2 , and $\operatorname{cos}\langle\cdot,\cdot\rangle$ the $\operatorname{cos}$ value of two vectors. The geometric regularization is defined as follows: $\mathcal{L}_{geo}=1-\frac{\operatorname{Cov}(\mathcal{R}_{D}(G_{t},\boldsymbol{0},\boldsymbol{d}_{i}),\Theta(\gamma(I,\boldsymbol{d}_{i})))}{\sqrt{\operatorname{Var}(\mathcal{R}_{D}(G_{t},\boldsymbol{0},\boldsymbol{d}_{i}))\operatorname{Var}(\Theta(\gamma(I,\boldsymbol{d}_{i})))}}$, where $\mathcal{R}_{D}$ is the rendered depth, $\operatorname{Cov}(\cdot,\cdot)$ the covariance, and $\operatorname{Var}(\cdot)$ the variance.
@@ 4 Experiments
Figure 5: Comparison between 4K4DGen and 3D-Cinemagraphy. 3D-Cinemagraphy produces ghost images at the middle frame.
@@ 4.1 Experimental Settings
We begin by detailing the acquisition of panoramas utilized in our experiments. These panoramas are generated from text prompts using state-of-the-art generative models feng2023diffusion360 ; yang2023dreamspace ; wu2023panodiffusion , initially at a resolution of $6144\times 3072$. Subsequently, these images are resized to $4096\times 2048$ or $2048\times 1024$. We further delineate the spherical representation details by selecting 20 uniform directions $\boldsymbol{u}$ on the sphere $\mathbb{S}^{2}$. In every experiment, the image plane size $\boldsymbol{s}$ is maintained at $0.6\times 0.6$, with a focal length $f=0.6$ and a resolution of $512\times 512$. Rotation along the z-axis is set to zero for all cameras, ensuring that the up-axis for the $i^{\rm{th}}$ camera aligns with the $(O,\boldsymbol{d}_{i},\boldsymbol{z})$ plane. During the animating phase, we employ the perspective denoiser $\Phi$, instantiated as the Animate-anything model dai2023animateanything , which fine-tunes the SVD model blattmann2023stable . In the Spatial-Temporal Geometric Alignment stage of the Lifting phase, the depth estimator $\Theta$ is represented by MiDaS ranftl2021vision ; birkl2023midas . Crucially, the depth map from the perspective image is adjusted by a scaling factor, corresponding to the projection of the unit-length ray direction onto the camera orientation $\boldsymbol{d}$. For scenes without boundaries, depth values for distant elements, such as the sky, are set to a finite number to facilitate optimization. The hyper-parameters are set as follows: $\lambda_{\rm{depth}}=1,\lambda_{\rm{scale}}=0.1,\lambda_{\rm{shift}}=0.01$. We conduct Spatial-Temporal Geometry Alignment optimization over 3000 iterations, with $\lambda_{\rm{scale}}$ and $\lambda_{\rm{shift}}$ set to zero during the first 1500 iterations. For the 4D representation training stage, Gaussian parameters are optimized over 10000 iterations for each time stamp $t$. The hyper-parameters for this stage are defined as $\lambda_{\rm{rgb}}=1,\lambda_{\rm{temporal}}=\lambda_{\rm{sem}}=\lambda_{\rm{geo}}=0.05$, and the disturbance vector range $\alpha$ is varied at $0.05$, $0.1$, and $0.2$ during the $5400$, $6600$, and $9000$ iterations, respectively. All experiments are executed on a single NVIDIA A100 GPU with 80 GB RAM.
@@
As there are no ground truth 4D scenes available, we render videos at specific test camera poses from the synthesized 4D representation and employ non-reference video/image quality assessment methods for quantitative evaluation of our approach. For the test views, we select random cameras with $\boldsymbol{p}=0$ as part of our testing camera set. We then introduce disturbances as described in Sec. 3.4, applying a disturbance factor of $\alpha=0.05$ at these selected views. Datasets. The task of generating 4D panoramas from static panoramas is novel, and thus, no pre-existing datasets are available. We assess our methodology using panoramas generated by text-to-panorama diffusion models, comprising a dataset of 16 panoramas. Baselines. Current SDS-based methods wu20234d ; zhao2023animate124 are limited to generating object-centered assets and do not support the generation of outward-facing scenes. We compare our method with the optical-flow-based 3D dynamic image technique, 3D-Cinemagraphy (3D-Cin.)li2023_3dcinemagraphy , by inputting the panorama into 3D-Cin. and projecting its output onto perspective images. Metrics. We assess video consistency using the CLIP consistency score, which averages the CLIP distances of consecutive frames, and the CLIP-I score, which evaluates the distance of each frame relative to the input panorama. Additionally, a user study is conducted where participants are solicited to vote for the video they perceive as superior in quality or decide whether two views are consistent with each other.
Table 1: Comparison with 3D-Cinemagraphy. Our method, 4K4DGen, consistently achieves higher CLIP similarity scores compared to the same panorama image processed by 3D-Cinemagraphy. Furthermore, the majority of participants in our user study rated the visual quality of 4K4DGen as superior.
Method CLIP(Consistency) $\uparrow$ CLIP-I $\uparrow$ User Vote $\uparrow$
3D-Cinemagraphy li2023_3dcinemagraphy 0.98 0.96 19%
4K4DGen 0.99 0.98 81%

### v3

@@ 3.2 Inconsistent Perspective Animation
Large-scale pre-trained 2D models have shown remarkable generative capabilities in creating images and videos, benefiting from vast multi-modal training data gathered from the Internet. However, acquiring high-quality 4D training data is considerably more challenging, and no current 4D dataset reaches the scale of those available for images and videos. Therefore, our approach aims to utilize the capabilities of video generative models to produce consistent panoramic 360 videos, which are then elevated to 4D. Nonetheless, the availability of panoramic videos is significantly more limited compared to planar perspective videos. Consequently, mainstream image-to-video (I2V) animation techniques may not perform optimally for panoramic formats, and the resolution of the videos remains constrained, as illustrated in Fig. 5 (b) and Tab. 2. Alternatively, the animator can be applied to perspective images. but this introduces inconsistencies across different projected views, as depicted in Fig. 5 (c)
@@ 3.3 Consistent Panoramic Animation
Limited by the scarcity of 4D training data in panoramic format, and given that large diffusion models are primarily trained on planar perspective videos, directly applying 2D perspective denoisers presents challenges in generating seamless panoramic videos with proper equirectangular projection, due to inconsistent motion across different views and the domain gap between spherical and perspective spaces. This constraint has driven us to develop a panoramic video generator in spherical space that leverages priors from general image-to-video (I2V) animation techniques, as shown in Fig. 2. Consequently, starting from a static input panorama, we animate it into a panoramic video, as demonstrated in the "Animating Phase" section of Fig. 3.
@@
To generate panoramic video from a static panorama, we build up the denoise-in-latent-space schema (An et al., 2023; Blattmann et al., 2023a; Dai et al., 2023) in a spherical context. For general video generation, a noisy latent sample is progressively denoised using DDPM  (Ho et al., 2020), conditioned on a static input image, and subsequently decoded into a video sequence by a pre-trained VAE decoder. However, in 4K4DGen, unlike the method for generating perspective planar videos, both the latent code and the static panorama input are represented on spheres. We start with the initial panoramic latent code $S^{T}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$, where $L$ denotes the number of video frames and $c$ the channels per frame. A novel Panoramic Denoiser is then applied to generate the clean panoramic latent code $S^{0}$, conditioned on the static input panorama $I\in{}^{H\times W}$. Subsequently, the equirectangular projection, as introduced in Sec. 3.1, projects the clean panoramic latent code into the matrix-like latent code $Z^{0}\in{}^{h\times w\times L\times c}$, with $h$ and $w$ representing the resolution of the latent code. Each $k^{\rm{th}}$ video frame $I^{k}$ in pixel space is decoded by the pre-trained VAE decoder as $I^{k}=\mathcal{D}(Z^{0}[:,:,k,:])$.
@@
We leverage a pre-trained perspective video generative model (Dai et al., 2023) to biuld our Panoramic Denoiser. This video generator takes a perspective image $i\in{}^{p_{H}\times p_{W}\times c}$ and an initial latent code $z^{T}\in{}^{p_{h}\times p_{w}\times(L\times c)}$ as inputs, progressively denoising the latent code $z^{T}$ to a clean state $z^{0}$ through a denoising function $z^{t-1}=\Phi(z^{t},i)$. Here, $p_{h}$ and $p_{w}$ represents the resolution of the latent code, $p_{H}$ and $p_{W}$ the resolution of the conditioning image, $c$ the number of channels, and $L$ the video length. Our goal is to transform the initial noisy panoramic latent code $S^{T}$ into the clean state $S^{0}$, ensuring that each perspective view is appropriately animated while maintaining global consistency. The underlying intuition is that if each perspective view undergoes its respective denoising process, the perspective video will feature meaningful animation. Moreover, if two perspective views overlap, they will align with each other  (Jiménez, 2023; Bar-Tal et al., 2023; Lugmayr et al., 2022) to produce a seamless global animation.
Given a static input panorama $I$ and an initial spherical latent code $S^{0}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$, we progressively remove noise employing a project-and-fuse procedure at each denoising step. Specifically, the spherical latent code at the $t^{\rm{th}}$ denoising step, $S^{t}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$, is projected into multiple perspective latent codes $\mathcal{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots,z^{t}_{n}\}$, where each $z^{t}_{k}=\gamma(S^{t},\bm{d}_{k})\in{}^{p_{h}\times p_{w}\times(L\times c)}$ represents the $k^{\rm{th}}$ perspective latent code projected in the equirectangular format detailed in Sec. 3.1. Each perspective latent code is then denoised by one step using a pre-trained perspective denoiser, denoted as $z_{k}^{t-1}=\Phi(z^{t}_{k},i_{k})$, where $i_{k}=\gamma(I,\bm{d}_{k})\in{}^{p_{H}\times p_{W}\times c}$ is the perspective conditioning image projected from the panorama $I$. Subsequently, we optimize the spherical latent code $S^{t-1}:\mathbb{S}^{2}\rightarrow{}^{L\times c}$ at step $t-1$ by fusing all the denoised perspective latent codes ${z_{k}^{t-1}}$. Formally, the denoising procedure at step $t$, denoted as $S^{t-1}=\Psi(S^{t},I)$, encompasses the following operations:
$\Psi\left(\mathcal{S}^{t},I\right)=\operatornamewithlimits{argmin}_{\mathcal{S}}\mathbb{E}_{\bm{d}\in\mathbb{S}^{2}}\|\gamma(\mathcal{S},\bm{d})-\Phi\left(\gamma(\mathcal{S}^{t},\bm{d}),\gamma(I,\bm{d})\right)\|.$ (3)
3.4 Dynamic Panoramic Lifting
@@
We define the panoramic video as $V=\{I^{1},I^{2},\ldots,I^{L}\}$, consisting of $L$ frames. The video is divided into overlapping perspective videos $\{v_{0},v_{1},\ldots,v_{n}\}$, each captured from specific camera directions $\{\bm{d}_{1},\ldots,\bm{d}_{n}\}$, collectively encompassing the entire span of the panoramic video $V$. Subsequently, we estimate the geometry of the 4D scene by fusing the depth maps through spatial-temporal geometry alignment. Following this, we describe our methodology for 4D representation and the subsequent rendering procedure.
@@
To transition from 2D video to 3D space, we utilize a monocular depth estimator  (Ranftl et al., 2021), inspired by advancements in  (Zhou et al., 2024b), to estimate the scene’s geometric structure. Nonetheless, depth maps generated for each frame and perspective might lack spatial and temporal consistency. To address this, we implement Spatial-Temporal Geometry Alignment using a pre-trained depth estimator $\Theta:{}^{h\times w\times 3}\rightarrow{}^{h\times w}$, applied to perspective images. Our objective is to amalgamate $n$ perspective depth maps $D^{K}_{i}=\Theta(\gamma(I^{k},\bm{d}_{i}))$ into a cohesive panoramic depth map $D^{k}$ for each frame $I^{k}$, ensuring spatial and temporal continuity. We express these depth maps as a spherical representation ${\mathcal{S}_{D}^{1},\ldots,\mathcal{S}_{D}^{L}}$. For enhanced optimization, we assign $n$ scale factors $\alpha_{i}^{k}\in\real$ and shifting parameters $\beta_{i}^{k}\in{}^{h\times w}$ to each perspective depth map. The comprehensive depth map $D^{k}$ is then optimized jointly with these parameters $\alpha$ and $\beta$. The formal objective is structured as follows:
@@
We represent and render the dynamic scene using $T$ sets of 3D Gaussians. Each set, corresponding to a specific timestamp $t$, is denoted as $G_{t}=\{\left(\bm{p}_{t}^{i},\bm{q}_{t}^{i},\bm{s}_{t}^{i},\bm{c}_{t}^{i},o_{t}^{i}\right)|i=1,\ldots,n\}$. This definition aligns with the methods described in  (Bahmani et al., 2023), which also provides a fast rasterizer for rendering images based on these Gaussian sets and given camera parameters. Consistent with Sec. 3.1, while the camera intrinsics remain fixed, we parameterize the camera extrinsics through a position $\bm{p}\in{}^{3}$ and an orientation $\bm{d}\in\mathbb{S}^{2}$. The training process is structured in two stages: initially, we directly supervise the 3D Gaussians using the panoramic videos. Let $\mathcal{R}(G,\bm{p},\bm{d})$ represent the rasterized image from Gaussian set $G$, utilizing camera extrinsics $\bm{p}=0$ and camera direction $\bm{d}$. Let $I_{t}$ denote the $t^{\rm{th}}$ frame of the panoramic video. We optimize the $t^{\rm{th}}$ Gaussian set $G_{t}$ using the following objective:
@@
where the RGB supervision term $\mathcal{L}_{\rm{rgb}}=\lambda\mathcal{L}_{1}+(1-\lambda)\mathcal{L}_{\rm{SSIM}}$ is the same as 3D-GS  (Kerbl et al., 2023), and the temporal regularize term $\mathcal{L}_{\rm{temporal}}$ written as:
$\mathcal{L}_{\rm{temporal}}=\sum\limits_{i=1}^{n}\|\mathcal{R}(G_{t},\bm{0},\bm{d}_{i})-\mathcal{R}(G_{t-1},\bm{0},\bm{d}_{i}))\|$ (6)
Then, we adopt the distillation loss and geometric regularization used in  (Zhou et al., 2024b), the distillation loss is defined as follows: $\mathcal{L}_{sem}=1-\operatorname{cos}\left\langle\operatorname{CLS}(\mathcal{R}(G_{t},\bm{0},\bm{d}_{i})),\operatorname{CLS}(\mathcal{R}(G_{t},\bm{\delta}_{p},\bm{d}_{i}))\right\rangle$, where $\bm{\delta}_{p}\in[-\alpha,\alpha]^{3}$ is the disturbing vector, $\operatorname{CLS}(\cdot)$ the feature extractor such as DINO (Oquab et al., 2023), and $\operatorname{cos}\langle\cdot,\cdot\rangle$ the $\operatorname{cos}$ value of two vectors. The geometric regularization is defined as follows: $\mathcal{L}_{geo}=1-\frac{\operatorname{Cov}(\mathcal{R}_{D}(G_{t},\bm{0},\bm{d}_{i}),\Theta(\gamma(I,\bm{d}_{i})))}{\sqrt{\operatorname{Var}(\mathcal{R}_{D}(G_{t},\bm{0},\bm{d}_{i}))\operatorname{Var}(\Theta(\gamma(I,\bm{d}_{i})))}}$, where $\mathcal{R}_{D}$ is the rendered depth, $\operatorname{Cov}(\cdot,\cdot)$ the covariance, and $\operatorname{Var}(\cdot)$ the variance.
@@ 4 Experiments
Figure 4: Comparison between 4K4DGen and 3D-Cinemagraphy. We present the input static panorama (Pano RGB), the corresponding text prompts, and the rendered results from different views and at various timestamps. 4K4DGen (Ours) effectively generates 4D scenes that are both spatially and temporally consistent, while 3D-Cinemagraphy (3D-Cin.) suffers from ghosting artifacts in the middle frames.
@@ 4.1 Experimental Settings
For perspective images, we uniformly select 20 directions $\bm{u}$ on the sphere $\mathbb{S}^{2}$ as the z-axis of 20 cameras. In each experiment, the image plane size $\bm{s}$ is set at $0.6\times 0.6$, with a focal length $f=0.6$ and a resolution of $512\times 512$. Rotation along the z-axis is kept at zero for all cameras, ensuring that the up-axis for the $i^{\rm{th}}$ camera aligns with the $(O,\bm{u}_{i},\bm{z})$ plane. During the animating phase, we utilize the perspective denoiser $\Phi$, instantiated as the Animate-anything model  (Dai et al., 2023), which fine-tunes the SVD model  (Blattmann et al., 2023a). In the Spatial-Temporal Geometric Alignment stage of the lifting phase, the depth estimator $\Theta$ is implemented using MiDaS  (Ranftl et al., 2021; Birkl et al., 2023). All experiments are executed on a single NVIDIA A100 GPU with 80 GB RAM.
@@
As there is no ground truth 4D scene data available, we render videos at specific test camera poses from the synthesized 4D representation and employ non-reference video/image quality assessment methods for quantitative evaluation of our approach. For the test views, we select random cameras with $\bm{p}=0$ as part of our testing camera set. We then introduce disturbances as described in Sec. 3.4, applying a disturbance factor of $\alpha=0.05$ at these selected views. Datasets. The task of generating 4D panoramas from static panoramas is new, and thus, no pre-existing datasets are available. In line with previous large-scale scene generation works (Zhou et al., 2024b; Yu et al., 2024), we evaluate our methodology using a dataset of 16 panoramas generated by text-to-panorama diffusion models. Baselines. Current SDS-based methods  (Wu et al., 2023a; Zhao et al., 2023) are limited to generating object-centered assets and do not support outward-facing scene generation. We compare our method with the optical-flow-based 3D dynamic image technique, 3D-Cinemagraphy (3D-Cin.) (Li et al., 2023b) (both the “circle” and “zoom-in” mode), by inputting the static panorama and projecting the output onto perspective images. Metrics. It is challenging to evaluate the visual quality without a ground-truth reference. We assess the rendered perspective videos based on both frame and video visual quality. For frame quality, we use distribution-based metrics such as FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018), which calculate the distance between generated frames and the corresponding perspective images projected from the static panoramas. We also employ the LLM-based visual scorer Q-Align (Wu et al., 2023b) to evaluate the quality of individual frames. For video quality, we use the Q-Align video model as the quality scorer. Additionally, we conduct user studies to further evaluate the results. In this paper, there are two types of user studies: (1) User Choice (UC), where participants are asked to compare and select the best video from candidates generated by different methods, and (2) User Agreement (UA), where participants assess whether specific properties are present in the videos generated by a particular approach.
Table 1: Comparison with 3D-Cinemagraphy. The IQ, IA, and VQ models represent the image quality scorer, image aesthetic scorer, and video quality scorer, respectively, within the Q-Align assessment framework. Our method, 4K4DGen, consistently achieves superior performance in both image and video quality across these metrics. Furthermore, the majority of participants in our user study rated 4K4DGen as the best in terms of visual quality.
Method FID $\downarrow$ KID (×$\times 10^{-2}$) $\downarrow$ Q-Align (IQ) $\uparrow$ Q-Align (IA) $\uparrow$ Q-Align (VQ) $\uparrow$ Quality (UC) $\uparrow$
3D-Cinemagraphy (zoom-in) 57.05 1.4 0.47 0.38 0.57 7%
3D-Cinemagraphy (circle) 57.78 1.3 0.48 0.40 0.58 12%
4K4DGen 16.59 0.2 0.66 0.44 0.62 81%